I have to make this media 5mM guanidine hcl YPD plate. the molecular weight for guanidine is 95.5 MW the YPD I make is 700ml. I need to make 100ml solution guanidine hcl first filter it before I add to the plate? Then I need to figure out how much guanidine hcl from that 100ml solution in mls I need to add to 700ml of the plates final conc 5mM


1
Expert's answer
2021-02-17T07:46:11-0500

m1v1= m2v2

m1 = m2v2/v1

m1 = 5×700/100 = 35 mM

no. Of moles = 3.5 moles

Mass required = 3.5×95.5

MASS = 334.25mg